OpenStudy (anonymous):

5 hockey pucks and 3 hockey sticks cost $23. 5 hockey pucks and and 1 hockey stick cost $20. How much does one hockey puck cost? Please show your work. Thank you in advance.

OpenStudy (campbell_st):

p = pucks and h = sticks 5p + 3h = 23 (1) 5p + h = 20 (2) solve simultaneously (1) -(2) 2h = 3 h = 1.50 sub back into (2) to find p 5p + 1.50 = 20 5p = 18.50 p = 3. 70 a hockey stick costs $1.50 a puck costs 3.70

OpenStudy (anonymous):

Thank you.